OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. g(x) = x/x^2-16

OpenStudy (whpalmer4):

Find the values of x that make the denominator equal to 0

OpenStudy (***[isuru]***):

check the denominator.. x^2 - 16 = ( x -4) (x + 4) any x value which will make ( x - 4) ( x+ 4) = 0 r the vertical asymptotes

OpenStudy (***[isuru]***):

can u figure them out now ?

OpenStudy (anonymous):

4,-4?

OpenStudy (anonymous):

ok I got it now, thank u sososo much!!!

OpenStudy (***[isuru]***):

yep..they r the answers

OpenStudy (***[isuru]***):

u r welcome :)